OpenStudy (anonymous):

I need to find sec theta...apparently my copy and paste didnt work so well...it is if sin theta 4/5 and tan theta is -4/3 , find sec theta

OpenStudy (freckles):

What is tan equal to in term of sin and cos

OpenStudy (anonymous):

tan is -4/3?? I typed out the entire question lol

OpenStudy (freckles):

yes but you didn't answer my question just now sin(t) /cos(t) = tan(t)

OpenStudy (freckles):

So what is tan(t) times 1/sin(t)

OpenStudy (anonymous):

well, 1/sin is how you find the csc, right?? Maybe?

OpenStudy (anonymous):

Idk, I'm trying to find an example in my book, and I can't

OpenStudy (freckles):

So you don't know that tan(t) =sin(t) /cos(t) and therefore tan(t) times 1/sin(t)=1/cos(t)=sec(t)

OpenStudy (anonymous):

well,yeah...you confused me

OpenStudy (freckles):

Hmm... What was confusing? And are you still confused? Do you think you can finish the question now?